Grade 12th passMechanics

A boy throws a ball in a train moving at a speed of 18 kmph with a velocity of 15√2 m/s at an angle of 45 to horizontal in the direction of motion of train. The angle with which it strikes the plane of projection w.r.t. an observer watching from ground near by to train

Mayank Vats
9 Years agoGrade 12th pass

2 Answers

Deepak Saxena
8 Years ago
-18 kmph=5 m/s
-v(x)=15+5=20{velocity along x before collision}
-v(y)=-15{velocity along y before collision is same as u(y) with opposite sign}
-now tan(Q)=v(y)/v(x)=15/20=3/4
so answer is 37 degrees
Basu
4 Years ago
-18 kmph=5 m/s
-v(x)=15+5=20{velocity along x before collision}
-v(y)=-15{velocity along y before collision is same as u(y) with opposite sign}
-now tan(Q)=v(y)/v(x)=15/20=3/4
so answer is 37 degree